# breaklines

### v3

@@ 10. Concluding Remarks
In this paper, we introduced AlphaTrans. This neuro-symbolic approach combines the power of static analysis and abilities of LLMs in code synthesis to automate repository-level code translation and validation. AlphaTrans decomposes the program into smaller fragments and translates the fragments in the reverse call order, originally building the project in the target language. In addition to syntax check, AlphaTrans implements two levels of validation through GraalVM and test translation. AlphaTrans is the first approach to translate and validate the entire repository, and we envision several research directions to advance repository-level code translation and validation.
One of the major challenges of repository-level code translation is identifying suitable library APIs in the target PL. Often, equivalent Python APIs may not exist, requiring new code generation or translation of the library API itself. Even if similar libraries exist, the logic of libraries might be different in two PLs. AlphaTrans supports translating frequently used APIs and aims to build a generic pipeline. Supporting all the libraries in the pipeline remains an open challenge we aim to address in future work. Furthermore, while the idea of compositional translation and validation is PL-agnostic, the static analysis makes the extension of AlphaTrans to translating from other source projects challenging. Devising LLM-enabled or PL-agnostic static analysis approaches can benefit code translation approaches such as AlphaTrans. We also showed that the quality of the source project test suite can significantly impact the translation validation results. As part of our future work, we plan to integrate an LLM-based test generator into the AlphaTrans pipeline to advance the validation component.
@@ 11. Data Availability
Artifacts and implementation of AlphaTrans are publicly available (Lab, 2025).

### v5

@@ 10. Concluding Remarks
In this paper, we introduced AlphaTrans, a neuro-symbolic approach that combines the power of static analysis and abilities of LLMs in code synthesis to automate repository-level code translation and validation. AlphaTrans decomposes the program into smaller fragments and translates the fragments in reverse call order, incrementally building the source project in the target language. In addition to syntactic checks, AlphaTrans implements two types of validation through GraalVM and test translation. AlphaTrans is the first approach to translate and validate real-world projects, and we envision several research directions to advance repository-level code translation and validation.
One of the major challenges in repository-level code translation is identifying suitable library APIs in the target PL. Often, equivalent Python APIs may not exist, requiring new code generation or translation of the library API itself. Even if similar libraries exist, the logic of libraries might be different in two PLs. AlphaTrans supports translating frequently used APIs and aims to build a generic pipeline. Supporting all the libraries in the pipeline remains an open challenge that we aim to address in future work. Furthermore, while the idea of compositional translation and validation is PL-agnostic, the static analysis makes the extension of AlphaTrans to translating from other source projects challenging. Devising LLM-enabled or PL-agnostic static analysis approaches can benefit code translation approaches such as AlphaTrans. We also showed that the quality of the source project test suite can significantly impact the translation validation results. In future work, we plan to integrate an LLM-based test generator into the AlphaTrans pipeline to enhance the validation component.
@@ 11. Data Availability
Artifacts and implementation of AlphaTrans are publicly available (Lab, 2025a, b).
Acknowledgements.
This work is supported by the IBM-Illinois Discovery Accelerator Institute and NSF CCF-2238045 grants. We thank Prof. Darko Marinov and Raju Pavuluri for their help with this research. We also thank the anonymous reviewers for their comments, which helped make this work stronger.